# A Comprehensive Analysis on LLM-based Node Classification Algorithms

### v1

@@ A.3 Zero-shot Learning with LLMs
On the other hand, GFMs are foundation models pre-trained on extensive graph corpora to achieve general graph intelligence. Approaches such as ZeroG (Li et al., 2024b) and OFA (Liu et al., 2024b) fine-tune LMs or GNNs on multiple graphs, enabling these models to generalize to unseen graph datasets without extensive retraining. There also exist other zero-shot learning methods utilizing LLMs. For example, Chen et al. (2024c) leverage LLMs as annotators to generate pseudo-labels for GNN training, enabling classification tasks. These approaches are not considered in this work, as our focus is primarily on LLMs directly solving node classification tasks in zero-shot scenarios.
@@ A.4 Benchmarks of LLMs for Graphs
The first category primarily utilizes LLMs for basic graph reasoning tasks, e.g., shortest path and connectivity. For instance, NLGraph (Wang et al., 2023) is a pioneering benchmark that encompasses eight different graph reasoning tasks presented in natural language. LLM4DyG (Zhang et al., 2024) further extends these reasoning tasks to dynamic graph settings. GraphArena (Tang et al., 2024) deals with more complex graph computational problems, with the complexity of tasks ranging from polynomial to NP-Complete like the Traveling Salesman Problem. ProGraph (Li et al., 2024a) evaluates the scalability of LLMs by handling large graphs with up to $10^{6}$ nodes, necessitating the use of Python APIs for graph analysis rather than relying solely on direct reasoning of LLMs. Additionally, Dai et al. (2024) investigates whether LLMs can recognize graph patterns, e.g., triangles or squares, based on terminological or topological descriptions.
The second category focuses on the potential of LLMs for node classification tasks. While numerous surveys discuss the progress in this area (Li et al., 2023; Jin et al., 2023a), benchmarks that systematically evaluate LLM-based node classification methods remain limited. In the preliminary work by Chen et al. (2024b), the exploration is confined to a narrow range of LLM-as-Encoder and LLM-as-Reasoner approaches, primarily focusing on a limited set of language models. GLBench (Li et al., 2024c) emerges as the first comprehensive benchmark for LLM-based node classification, offering consistent data splits to evaluate representative methods in both semi-supervised and zero-shot settings. However, variations in backbone models and implemented codebases impede fair and rigorous comparisons.
@@ C.1 Datasets
Table 8: Statistics of supported datasets in LLMNodeBed.
Academic Web Link Social E-Commerce
Statistics Cora Citeseer Pubmed arXiv WikiCS Instagram Reddit Books Photo Computer
# Classes 7 6 3 40 10 2 2 12 12 10
# Nodes 2,708 3,186 19,717 169,343 11,701 11,339 33,434 41,551 48,362 87,229
# Edges 5,429 4,277 44,338 1,166,243 215,863 144,010 198,448 358,574 500,928 721,081
Avg. # Token 183.4 210.0 446.5 239.8 629.9 56.2 197.3 337.0 201.5 123.1
Homophily (%) 82.52 72.93 79.24 63.53 68.67 63.35 55.52 78.05 78.50 85.28
@@
Setting Cora Citeseer Pubmed arXiv WikiCS Instagram Reddit Books Photo Computer
Semi-supervised 5.17% 3.77% 0.30% - 4.96% 10.00% 10.00% 10.00% 10.00% 10.00%
Supervised 60.0% 60.0% 60.0% 53.7% 60.0% 60.0% 60.0% 60.0% 60.0% 60.0%
We selected 10 datasets from academic, web link, social, and e-commerce domains to create a diverse graph database. Within LLMNodeBed, each dataset is stored in .pt format using PyTorch, which includes shallow embeddings, raw text of nodes, edge indices, labels, and data splits for convenient loading. The datasets are described below, while their statistics and additional details are provided in Table 8 and Table 9, respectively.
  • Academic Networks: The Cora (Sen et al., 2008), Citeseer (Giles et al., 1998), Pubmed (Yang et al., 2016b), and ogbn-arXiv (abbreviated as “arXiv”) (Hu et al., 2020) datasets consist of nodes representing papers, with edges indicating citation relationships. The associated text attributes include each paper’s title and abstract, which we use the collected version as follows: Cora and Pubmed from He et al. (2023), Citeseer from Chen et al. (2024b). Within the dataset, each node is labeled according to its category. For example, the arXiv dataset includes 40 CS sub-categories such as cs.AI (Artificial Intelligence) and cs.DB (Databases).
  • Web Link Network: In the WikiCS dataset (Mernyei & Cangea, 2020), each node represents a Wikipedia page, and edges indicate reference links between pages. The raw text for each node includes the page name and content, which was collected by Liu et al. (2024b). The classification goal is to categorize each entity into different Wikipedia categories.
@@ C.2 Implementation Details and Hyperparameters Setting
    For both GNNs and MLPs across experimental datasets, the learning_rate is consistently set to $21e-$ 2, following previous studies (He et al., 2023; Li et al., 2024c). The total number of epochs is set to $500$ with a patience of $100$.
@@
  • For ENGINE (Zhu et al., 2024b), we refer to the hyperparameter settings outlined in the original paper to determine the hyperparameter search space as follows:
@@
  • For TAPE (He et al., 2023), we utilize the provided prompt templates to guide Mistral-7B and GPT-4o in conducting reasoning. The LM, RoBERTa-355M, is fine-tuned based on its default parameter settings, while the GNN hyperparameters are explored with num_layers in $[2,3,4]$, hidden_dimension in $[128,256]$, and with or without batch_normalization.
  • For LLM Instruction Tuning, we use the LoRA (Hu et al., 2021) techniques to fine-tune LLMs. The lora_r parameter (dimension for LoRA update matrices) is set to $8$ and the lora_alpha (scaling factor) to $16$. The dropout ratio is set to $0.1$, the batch_size to $16$, and the learning_rate to $51e-$ 5. For each dataset, the input consists of the node’s original text along with a carefully crafted task prompt designed to guide the LLMs in performing the classification task. The expected output is the corresponding label. For small-scale datasets such as Cora, Citeseer, and Instagram, the number of training epochs is $10$ in semi-supervised settings and $4$ in supervised settings. For the remaining datasets, the training epochs are $2$ and $1$ for semi-supervised and supervised settings, respectively. The maximum input and output lengths are determined based on the average token lengths of each dataset.
@@ C.3 Distinct Features
    • GNNs: Our framework supports a wide range of variants, including GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), GIN (Xu et al., 2019), and Graph Transformers (Shi et al., 2020). These GNNs can be customized with various layers and embedding dimensions.
    • LMs and LLMs: Open-source models can be easily loaded via the Transformers library [^2] https://huggingface.co/docs/transformers. In our experiments, we primarily utilize SenBERT-66M (Reimers & Gurevych, 2019), RoBERTa-355M (Liu et al., 2019), Qwen2.5-Series (Yang et al., 2024), Mistral-7B (Jiang et al., 2023), and LLaMA3.1-8B (Dubey et al., 2024). For close-source LLMs, we have formatted the invocation functions of DeepSeek-Chat (Liu et al., 2024a) and GPT-4o (OpenAI, 2024). Additionally, LLMNodeBed allows users to specify and invoke any LM or LLM of their choice, providing flexibility for diverse research needs.
@@
Appendix D Supplementary Materials for Comparisons among Algorithm Categories
@@
Reasoner TAPE 81.89±${}_{\pm\text{0.31}}$ 66.80±${}_{\pm\text{0.33}}$ 78.46±${}_{\pm\text{1.13}}$ 80.03±${}_{\pm\text{0.23}}$ 50.01±${}_{\pm\text{1.60}}$ 61.23±${}_{\pm\text{0.69}}$ 47.12±${}_{\pm\text{3.26}}$ 82.31±${}_{\pm\text{0.19}}$ 84.90±${}_{\pm\text{1.14}}$ 70.31
@@
Reasoner TAPE 87.21±${}_{\pm\text{1.60}}$ 73.33±${}_{\pm\text{1.57}}$ 92.39±${}_{\pm\text{0.02}}$ 57.79±${}_{\pm\text{0.49}}$ 86.03±${}_{\pm\text{1.14}}$ 58.31±${}_{\pm\text{1.15}}$ 65.91±${}_{\pm\text{0.71}}$ 54.07±${}_{\pm\text{2.01}}$ 83.41±${}_{\pm\text{0.42}}$ 86.78±${}_{\pm\text{0.53}}$ 74.52
@@
Table 11: Performance comparison under zero-shot setting with both Accuracy (%) and Macro-F1 (%) reported.
@@
Appendix E Supplementary Materials for Fine-grained Analysis
E.1 LLM-as-Encoder: Compared with LMs
Table 12: Comparison of LLM- and LM-as-Encoder with Accuracy (%) reported under supervised setting. LLM-as-Encoder outperforms LMs in heterophilic graphs. The best encoder within each method on a dataset is highlighted.
@@ E.1 LLM-as-Encoder: Compared with LMs
We supplement the comparison between LLM-as-Encoder and LM-as-Encoder under supervised settings in Table 12. The key takeaway that LLMs outperform LMs as Encoders in heterophilic graphs remains valid. This conclusion is particularly evident on the arXiv dataset, where the performance gap between LM- and LLM-generated embeddings reaches up to $7\%$ on MLP and $3\%$ on GraphSAGE. Additionally, we observe that in supervised settings, the performance gap between LM- and LLM-as-Encoders becomes less pronounced compared to semi-supervised settings. We still consider the theoretical insights in Equation (1) for explanation: Increased supervision enhances the mutual information between labels and graph structure, i.e., $I(\mathcal{E},\mathcal{Y}_{l})$, thereby rendering the second term less significant and diminishing the advantages provided by more powerful encoders like LLMs.
E.2 LLM-as-Predictor: Sensitivity to LLM Backbones
@@
E.3 LLM-as-Predictor: Biased and Hallucinated Predictions
@@
Appendix F Supplementary Materials for Computational Cost Analysis
Table 17: Total training times of different methods in semi-supervised settings. All recorded experiment times are based on a single NVIDIA H100-80G GPU.
Type Method Cora Citeseer Pubmed WikiCS Instagram Reddit Books Photo Computer
# Training Samples 140 120 60 580 1,160 3,344 4,155 4,836 8,722
[…]
SenBERT-66M 8.5s 7.9s 5.9s 27.9s 14.7s 1.2m 1.5m 1.8m 3.3m
RoBERTa-355M 21.2s 18.9s 12.7s 1.2m 2.3m 6.5m 8.1m 3.8m 6.9m
[…]
ENGINE 2.2m 2.4m 16.1m 15.2m 9.3m 22.9m 31.1m 38.8m 1.1h
Reasoner TAPE 25.5m 27.8m 5.6h 2.7h 2.0h 8.0h 9.9h 11.7h 14.5h
[…]
GraphGPT 16.4m 15.5m 1.2h 48.5m 30.1m 1.8h 2.4h 2.2h 5.8h
LLaGA 1.7m 2.2m 5.2m 5.8m 3.0m 20.9m 19.5m 23.5m 43.1m

### v2

@@ A.3 Zero-shot Learning with LLMs
On the other hand, GFMs are foundation models pre-trained on extensive graph corpora to achieve general graph intelligence. Approaches such as ZeroG (Li et al., 2024b) and OFA (Liu et al., 2024) fine-tune LMs or GNNs on multiple graphs, enabling these models to generalize to unseen graph datasets without extensive retraining. There also exist other zero-shot learning methods utilizing LLMs. For example, Chen et al. (2024c) leverage LLMs as annotators to generate pseudo-labels for GNN training, enabling classification tasks. These approaches are not considered in this work, as our focus is primarily on LLMs directly solving node classification tasks in zero-shot scenarios.
@@ A.4 Benchmarks of LLMs for Graphs
The first category primarily utilizes LLMs for basic graph reasoning tasks, e.g., shortest path and connectivity. For instance, NLGraph (Wang et al., 2023) is a pioneering benchmark that encompasses eight different graph reasoning tasks presented in natural language. LLM4DyG (Zhang et al., 2024) further extends these reasoning tasks to dynamic graph settings. GraphArena (Tang et al., 2025) deals with more complex graph computational problems, with the complexity of tasks ranging from polynomial to NP-Complete like the Traveling Salesman Problem. ProGraph (Li et al., 2024a) evaluates the scalability of LLMs by handling large graphs with up to $10^{6}$ nodes, necessitating the use of Python APIs for graph analysis rather than relying solely on direct reasoning of LLMs. Additionally, Dai et al. (2025) investigates whether LLMs can recognize graph patterns, e.g., triangles or squares, based on terminological or topological descriptions.
The second category focuses on the potential of LLMs for node classification tasks. While numerous surveys discuss the progress in this area (Li et al., 2023b; Jin et al., 2023a), benchmarks that systematically evaluate LLM-based node classification methods remain limited. In the preliminary work by Chen et al. (2024b), the exploration is confined to a narrow range of LLM-as-Encoder and LLM-as-Explainer approaches, primarily focusing on a limited set of language models. GLBench (Li et al., 2024c) emerges as the first comprehensive benchmark for LLM-based node classification, offering consistent data splits to evaluate representative methods in both semi-supervised and zero-shot settings. However, variations in backbone models and implemented codebases impede fair and rigorous comparisons.
@@ C.1 Datasets
Heterophilic Student, Faculty, Staff, Course, Project
Setting Cora Citeseer Pubmed arXiv WikiCS Instagram Reddit Books Photo Computer Heterophilic
Semi-supervised 5.17% 3.77% 0.30% - 4.96% 10.00% 10.00% 10.00% 10.00% 10.00% 10.00%
Supervised 60.0% 60.0% 60.0% 53.7% 60.0% 60.0% 60.0% 60.0% 60.0% 60.0% 60.0%
We selected 14 datasets from academic, web link, social, and e-commerce domains to create a diverse graph database. Within LLMNodeBed, each dataset is stored in .pt format using PyTorch, which includes shallow embeddings, raw text of nodes, edge indices, labels, and data splits for convenient loading. The processed data is publicly available at https://huggingface.co/datasets/xxwu/LLMNodeBed. A description of the datasets is provided below, with their statistics and additional details summarized in Table 1 and Table 10, respectively.
  • Academic Networks: The Cora (Sen et al., 2008), Citeseer (Giles et al., 1998), Pubmed (Yang et al., 2016), and ogbn-arXiv (abbreviated as “arXiv”) (Hu et al., 2020) datasets consist of nodes representing papers, with edges indicating citation relationships. The associated text attributes include each paper’s title and abstract, which we use the collected version as follows: Cora and Pubmed from He et al. (2024), Citeseer from Chen et al. (2024b). Within the dataset, each node is labeled according to its category. For example, the arXiv dataset includes 40 CS sub-categories such as cs.AI (Artificial Intelligence) and cs.DB (Databases).
  • Web Link Network: In the WikiCS dataset (Mernyei & Cangea, 2020), each node represents a Wikipedia page, and edges indicate reference links between pages. The raw text for each node includes the page name and content, which was collected by Liu et al. (2024). The classification goal is to categorize each entity into different Wikipedia categories.
@@
  • Heterophilic Datasets: The Cornell, Texas, Wisconsin, and Washington datasets are collected from Wang et al. (2025), which are the only available heterophilic datasets that have text attributes. Since the original release does not provide shallow embeddings, we generate fixed 300-dimensional embeddings for each node using Node2Vec (Grover & Leskovec, 2016). For dataset splits, we assign semi-supervised and supervised settings with 1:1:8 and 6:2:2 splits for training, validation, and test sets, respectively.
@@ C.2 Implementation Details and Hyperparameters Setting
    For the heterophilic GNN, H2GCN (Zhu et al., 2020), which we employed on heterophilic graphs in the fine-grained analysis of LLM-as-Encoder, we perform a grid-search over the following hyperparameters:
    num_layers in $[1,2,3]$, hidden_dimension in $[32,64,128,256]$, and learning_rate in $[0.001,0.005,0.01]$.
@@
    For both GNNs and MLPs across experimental datasets, the learning_rate is consistently set to $21e-$ 2, following previous studies (He et al., 2024; Li et al., 2024c). The total number of epochs is set to $500$ with a patience of $100$.
@@
  • For GLEM (Zhao et al., 2023a), we uniformly set the number of EM iterations to $1$ and the pseudo-labeling ratio to $0.5$. For the GNN module within GLEM, we set the hidden_dim to $256$ and num_layers to $3$. For the LM module, we use LoRA for optimization with a batch_size of $32$. Except for the Pubmed, arXiv, and Computer datasets, where the LM is trained first during the EM iteration, all remaining datasets train the GNN first. This choice is based on empirical findings that suggest better performance.
  • For ENGINE (Zhu et al., 2024), we refer to the hyperparameter settings outlined in the original paper to determine the hyperparameter search space as follows:
@@
  • For TAPE (He et al., 2024), we utilize the provided prompt templates to guide Mistral-7B and GPT-4o in conducting reasoning. The LM, RoBERTa-355M, is fine-tuned based on its default parameter settings, while the GNN hyperparameters are explored with num_layers in $[2,3,4]$, hidden_dimension in $[128,256]$, and with or without batch_normalization.
  • For LLM Instruction Tuning, we use the LoRA (Hu et al., 2022) techniques to fine-tune LLMs. The lora_r parameter (dimension for LoRA update matrices) is set to $8$ and the lora_alpha (scaling factor) to $16$. The dropout ratio is set to $0.1$, the batch_size to $16$, and the learning_rate to $51e-$ 5. For each dataset, the input consists of the node’s original text along with a carefully crafted task prompt designed to guide the LLMs in performing the classification task. The expected output is the corresponding label. For small-scale datasets such as Cora, Citeseer, and Instagram, the number of training epochs is $10$ in semi-supervised settings and $4$ in supervised settings. For the remaining datasets, the training epochs are $2$ and $1$ for semi-supervised and supervised settings, respectively. The maximum input and output lengths are determined based on the average token lengths of each dataset.
@@ C.3 Distinct Features
    • GNNs: Our framework supports a wide range of variants, including GCN (Kipf & Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), GIN (Xu et al., 2019), Graph Transformers (Shi et al., 2021) and H2GCN (Zhu et al., 2020). These GNNs can be customized with various layers and embedding dimensions.
    • LMs and LLMs: Open-source models can be easily loaded via the Transformers library [^3] https://huggingface.co/docs/transformers. In our experiments, we primarily utilize SenBERT-66M (Reimers & Gurevych, 2019), RoBERTa-355M (Liu et al., 2019), Qwen2.5-Series (Qwen et al., 2025), Mistral-7B (Jiang et al., 2023), and LLaMA3.1-8B (Dubey et al., 2024). For close-source LLMs, we have formatted the invocation functions of DeepSeek-V3 (DeepSeek-AI, 2024) and GPT-4o (OpenAI, 2024). Additionally, LLMNodeBed allows users to specify and invoke any LM or LLM of their choice, providing flexibility for diverse research needs.
@@
Appendix D Supplementary Experiments on Large-scale Dataset
[…]
Classic Encoder Predictor
Metric GCNShallowEmb SAGEShallowEmb GATShallowEmb SenBERT RoBERTa GCNLLMEmb SAGELLMEmb GATLLMEmb LLaGA
Accuracy 75.16 77.33 74.83 69.06 75.12 81.09 83.34 82.20 79.07
Macro-F1 36.16 35.63 34.32 24.89 30.29 39.96 41.14 40.49 36.36
We extend our evaluation to the large-scale graph, ogbn-products (Hu et al., 2020), which comprises 2,449,029 nodes and 123,718,152 edges. In this graph, each node represents a unique product sold on Amazon, and edges denote co-purchase relationships. The classification task involves predicting the category of a product across 47 classes. This dataset is significantly larger than any other in LLMNodeBed, posing unique challenges for evaluating LLM-based node classification algorithms.
Experimental Setups: To evaluate LLM-based methods under constrained GPU resources (e.g., a single H100-80GB GPU), we use the official data splits, which include 196,165 nodes (8.03%) for training, 39,323 nodes (1.61%) for validation, and a reduced test set of 221,309 nodes (9.04%), yielding a semi-supervised setting. The original dataset provides 100-dimensional node embeddings, which we directly adopt as shallow embeddings for classic GNNs.
Compared Methods: Given resource constraints, we compare classic methods (GNNs and LMs), LLM-as-Encoder methods (using embeddings derived from Qwen2.5-3B), and the LLM-as-Predictor method, LLaGA. Some other methods, such as TAPE, are excluded due to the prohibitive cost of generating explanatory texts for the entire dataset via APIs, e.g., approximately 500 USD for invoking GPT-4o.
[…]
Appendix E Supplementary Materials for Comparisons among Algorithm Categories
@@
GLEM 79.82±${}_{\pm\text{0.95}}$ 64.66±${}_{\pm\text{2.39}}$ 81.92±${}_{\pm\text{0.97}}$ 73.88±${}_{\pm\text{0.25}}$ 52.09±${}_{\pm\text{5.81}}$ 49.30±${}_{\pm\text{4.48}}$ 43.33±${}_{\pm\text{3.57}}$ 67.53±${}_{\pm\text{1.18}}$ 66.01±${}_{\pm\text{1.33}}$ 64.28
@@
Explainer TAPE 81.89±${}_{\pm\text{0.31}}$ 66.80±${}_{\pm\text{0.33}}$ 78.46±${}_{\pm\text{1.13}}$ 80.03±${}_{\pm\text{0.23}}$ 50.01±${}_{\pm\text{1.60}}$ 61.23±${}_{\pm\text{0.69}}$ 47.12±${}_{\pm\text{3.26}}$ 82.31±${}_{\pm\text{0.19}}$ 84.90±${}_{\pm\text{1.14}}$ 70.31
@@
GLEM 85.89±${}_{\pm\text{0.72}}$ 70.06±${}_{\pm\text{2.02}}$ 93.43±${}_{\pm\text{0.35}}$ 57.99±${}_{\pm\text{1.29}}$ 77.60±${}_{\pm\text{0.86}}$ 58.87±${}_{\pm\text{2.79}}$ 45.05±${}_{\pm\text{7.72}}$ 46.96±${}_{\pm\text{3.55}}$ 69.60±${}_{\pm\text{1.77}}$ 68.62±${}_{\pm\text{1.14}}$ 67.41
@@
Explainer TAPE 87.21±${}_{\pm\text{1.60}}$ 73.33±${}_{\pm\text{1.57}}$ 92.39±${}_{\pm\text{0.02}}$ 57.79±${}_{\pm\text{0.49}}$ 86.03±${}_{\pm\text{1.14}}$ 58.31±${}_{\pm\text{1.15}}$ 65.91±${}_{\pm\text{0.71}}$ 54.07±${}_{\pm\text{2.01}}$ 83.41±${}_{\pm\text{0.42}}$ 86.78±${}_{\pm\text{0.53}}$ 74.52
@@
[…]
@@
Appendix F Supplementary Materials for Fine-grained Analysis
F.1 LLM-as-Encoder: Compared with LMs
[…]
@@ F.1 LLM-as-Encoder: Compared with LMs
We supplement the comparison between LLM-as-Encoder and LM-as-Encoder under supervised settings in Table 14. The key takeaway that LLMs outperform LMs as Encoders in less informative graphs, e.g., heterophilic graphs remains valid. This conclusion is particularly evident on the arXiv dataset, where the performance gap between LM- and LLM-generated embeddings reaches up to $7\%$ on MLP and $3\%$ on GraphSAGE. Additionally, we observe that in supervised settings, the performance gap between LM- and LLM-as-Encoders becomes less pronounced compared to semi-supervised settings. We still consider the theoretical insights in Equation (1) for explanation: Increased supervision enhances the mutual information between labels and graph structure, i.e., $I(\mathcal{E},\mathcal{Y}_{l})$, thereby rendering the second term less significant and diminishing the advantages provided by more powerful encoders like LLMs.
F.2 LLM-as-Predictor: Sensitivity to LLM Backbones
@@
F.3 LLM-as-Predictor: Biased and Hallucinated Predictions
@@
Appendix G Supplementary Materials for Computational Cost Analysis